# Beyond Cosine Similarity: Magnitude-Aware CLIP for No-Reference Image Quality Assessment

### v1

@@ Introduction
Early no-reference IQA approaches relied on handcrafted features that captured natural scene statistics (NSS), with descriptors derived from spatial (mittal2012making), wavelet (moorthy2010two), and DCT (saad2012blind) domains. Psychovisual models inspired by the free-energy principle of the human visual system further enriched the modeling of perceptual degradation (gu2014using; zhai2011psychovisual). With the advent of deep learning, CNN-based models became prevalent. Early works such as Kang et al. (kang2014convolutional) learned quality-aware representations directly from image patches, and subsequent extensions introduced multi-task training (kang2015simultaneous) to jointly predict quality and distortion types. Later models such as DIQaM (bosse2017deep), RankIQA (liu2017rankiqa), FPR (chen2020no), and GraphIQA (sun2022graphiqa) further advanced performance by modeling spatial dependencies and incorporating relational reasoning. Transformer-based models have also been explored due to their strong modeling capacity and flexibility. You et al. (you2021transformer) introduced a Transformer backbone for NR-IQA that benefits from features pretrained on large-scale classification tasks. Other approaches leverage perceptual priors from image restoration networks (lin2018hallucinated; chen2022no). However, these models struggle to generalize due to limited data and domain-specific bias. To address this, recent work incorporates auxiliary tasks (you2021transformer; lin2018hallucinated) and learning strategies such as meta-learning (zhu2020metaiqa), curriculum learning (wang2023hierarchical), and domain generalization (chen2021learning; chen2025monotonic). Despite progress, generalization in NR-IQA remains an open problem.
@@
Figure 3: Semantic Bias exists in CLIP feature. This feature magnitudes for visually similar-quality images differ substantially across semantic categories. Statistical normalization is necessary to make magnitude cues reliable. WD represents the Wasserstein Distance between two Feature distribution.

### v2

@@ Introduction
Early no-reference IQA approaches relied on handcrafted features that captured natural scene statistics (NSS), with descriptors derived from spatial (mittal2012making), wavelet (moorthy2010two), and DCT (saad2012blind) domains. Psychovisual models inspired by the free-energy principle of the human visual system further enriched the modeling of perceptual degradation (gu2014using; zhai2011psychovisual). With the advent of deep learning, CNN-based models became prevalent. Early works such as Kang et al. (kang2014convolutional) learned quality-aware representations directly from image patches, and subsequent extensions introduced multi-task training (kang2015simultaneous) to jointly predict quality and distortion types. Later models such as DIQaM (bosse2017deep), FPR (chen2020no), and GraphIQA (sun2022graphiqa) further advanced performance by modeling spatial dependencies and incorporating relational reasoning. Transformer-based models have also been explored due to their strong modeling capacity and flexibility. You et al. (you2021transformer) introduced a Transformer backbone for NR-IQA that benefits from features pretrained on large-scale classification tasks. Other approaches leverage perceptual priors from image restoration networks (lin2018hallucinated; chen2022no). However, these models struggle to generalize due to limited data and domain-specific bias. To address this, recent work incorporates auxiliary tasks (you2021transformer; lin2018hallucinated) and learning strategies such as meta-learning (zhu2020metaiqa), curriculum learning (wang2023hierarchical), and domain generalization (chen2021learning; chen2025monotonic). Despite progress, generalization in NR-IQA remains an open problem.
@@
Figure 3: Semantic Bias exists in CLIP feature. This feature magnitudes for visually similar-quality images differ substantially across semantic categories. Statistical normalization is vital to make magnitude cues reliable. WD represents the Wasserstein Distance between two Feature distribution.